An electron (-q= +1.602 × 10-¹9) moves along a straight line from point b with a separation distance of 0.2 m.Considering the electric field along this line is uniform with magnitude of E=300 N/C and directed from points a to b.What is the magnitude of force electron?


1
Expert's answer
2022-03-06T15:18:03-0500

Solution

Force on electron

F=qE

"=1.602*10^{-19}*300\\=4.81*10^{-19} J"